# Edge of Many-Body Quantum Chaos in Quantum Reservoir Computing

### v1

@@ IV Reservoir performance
The main text presents QRC performance on the short-term memory (STM) task and the nonlinear auto-regressive moving average (NARMA) task. Specifically, it benchmarks the SYK4 [Fig. 3(c)] and SYK2 [Fig. 3(d)] models in the temporal domain, as well as the interpolated model in the parametric domain at input intervals of $\Delta t_{\mathrm{in}}=50$ [Fig. 4(a)] and $\Delta t_{\mathrm{in}}=1$ [Fig. 4(b)]. The aim here is to provide additional data to further elucidate QRC performance and the edge of many-body quantum chaos.
@@ IV.1 Reservoir performance for the SYK4 and SYK2 models
The tendency of short input intervals to introduce atypical decay in $R^{2}_{d}$ is more pronounced in the SYK2 model [Fig. S8(b)]. For instance, at $\Delta t_{\mathrm{in}}\leq 1$, the memory performance drops abruptly instead of decaying gradually as in the Haar QRC reference. This behavior suggests a compromised fading-memory property; due to the slow convergence, the system fails to sufficiently prioritize recent information. Nevertheless, for longer input intervals ($\Delta t_{\mathrm{in}}=5$, $10$, and $500$), QRC with the SYK2 model demonstrates significantly better memory retention than the Haar QRC baseline, with a long tail spanning beyond $d=50$. Importantly, this enhancement appears even though the convergence speed [quantified by $||\rho_{A}^{(k)}-\rho_{B}^{(k)}||$ in Fig. S6(b)] is comparable to that observed at shorter $\Delta t_{\mathrm{in}}$. This discrepancy implies that, while the distance-based convergence metric is valuable for characterizing QRC systems, its direct connection to performance can be system-dependent. Specifically, in the SYK2 model, ESP is satisfied only asymptotically, and it remains nontrivial to determine how this asymptotic nature precisely affects performance. Further investigation, potentially employing alternative convergence metrics, is left for future work.
@@
Figures S8(c) and S8(d) extend Figs. S8(a) and S8(b) respectively, by presenting heatmaps of $R_{d}^{2}$ for $0\leq d\leq 20$ as the input interval $\Delta t_{\mathrm{in}}$ is varied. In the SYK4 model [Fig. S8(c)], a key observation is an oblique line of high performance that appears where $\Delta t_{\mathrm{in}}$ is shorter than the Thouless time. This sharp feature represents the performance enhancement at the temporal edge of many-body quantum chaos, consistent with the observation in Fig. 3(c) of the main text. Since the Thouless time marks the point at which the QRC performance converges to that of the Haar QRC model, any enhancement should occur for $\Delta t_{\mathrm{in}}$ below this time. The exact location of the peak is task-dependent; for the task investigated here, it weakly depends on $d$, leading to the observed oblique line. Remarkably, a similar enhancement is absent in the SYK2 model [Fig. S8(d)]; instead, it shows a binary separation between a regime of sufficient memory retention at large $\Delta t_{\mathrm{in}}$ and poor retention at small $\Delta t_{\mathrm{in}}$, lacking any distinct performance peak.
@@
Figures S8(e) and S8(f) show heatmaps of the NMSE for the NARMA tasks using the SYK4 and SYK2 models, respectively. The vertical axis shows the NARMA order $n$, and the horizontal axis indicates the input interval $\Delta t_{\mathrm{in}}$. Consistent with the STM task results, a distinct performance peak is evident in the SYK4 model, whereas the SYK2 model shows only gradual performance changes. As discussed in the main text, the absence of peak-like structures in the SYK2 model supports the interpretation that the performance enhancement observed in the SYK4 model indeed originates from many-body quantum chaos. This enhancement persists for larger STM delays $d$ and NARMA orders $n$; although the peak broadens as task complexity increases, the overall observation confirms the robustness of the performance boost induced by the temporal edge.
@@ IV.2 Reservoir performance for the interpolated model
We additionally present performance heatmaps for the NARMA task in Fig. S9(e) for $\Delta t_{\mathrm{in}}=50$ and Fig. S9(f) for $\Delta t_{\mathrm{in}}=1$. The horizontal axis represents the coupling strength $\kappa_{2}/J_{4}$, and the vertical axis shows the NARMA order $n$. For small NARMA orders with a long input interval [Fig. S9(e)], the performance is distinctly enhanced around the chaotic boundary, corresponding to the parametric edge of many-body quantum chaos. This edge-induced performance enhancement is absent for the short input interval $\Delta t_{\mathrm{in}}=1$ in Fig. S9(f) because, as discussed in the main text, the QRC requires a sufficiently long input interval to manifest its quantum chaotic characteristics.
A comparison between the STM task [Fig. S9(c)] and the NARMA task [Fig. S9(e)] illuminates distinct aspects of performance enhancement. Specifically, for small values of $d$ and $n$, the NARMA performance displays a pronounced peak around the chaotic boundary, whereas the STM performance shows no comparable improvement. Given that the NARMA task requires both memory and nonlinear processing, we conclude that enhancements in nonlinear transformation, rather than memory, primarily drive the observed performance boost at this parametric edge. As the NARMA order $n$ increases, however, the task requires more memory, and memory capacity becomes the dominant performance factor. Consequently, the NARMA performance displays an oblique contour in the high-$n$ region, similar to the STM performance, and the edge-induced peak is gradually obscured.

### v3

@@ IV Reservoir performance
The main text presents QRC performance on the short-term memory (STM) task and the nonlinear auto-regressive moving average (NARMA) task. Specifically, it benchmarks the SYK4 [Fig. 3(c)] and SYK2 [Fig. 3(d)] models in the temporal domain, as well as the interpolated model in the parametric domain at input intervals of $\Delta t_{\mathrm{in}}=50$ [Fig. 4(a)] and $\Delta t_{\mathrm{in}}=1$ [Fig. 4(b)]. The aim here is to provide additional data to further elucidate QRC performance and the edge of many-body quantum chaos. In Secs. IV.1 and IV.2, we examine the STM and NARMA performance in the temporal and parametric domains, respectively, by sweeping the delay parameters and task orders. We also report the QRC performance on additional tasks: the nonlinear STM task in Sec. IV.3 and the parity-check (PC) task in Sec. IV.4. Finally, Sec. IV.5 investigates the effects of statistical noise on QRC performance.
@@ IV.1 Reservoir performance for the SYK4 and SYK2 models
The tendency of short input intervals to introduce atypical decay in $R^{2}_{d}$ is more pronounced in the SYK2 model [Fig. S8(b)]. For instance, at $\Delta t_{\mathrm{in}}\leq 1$, the memory performance drops abruptly instead of decaying gradually. This behavior suggests a compromised fading-memory property; due to the slow convergence, the system fails to sufficiently prioritize recent information. Nevertheless, for longer input intervals ($\Delta t_{\mathrm{in}}=5$, $10$, and $500$), QRC with the SYK2 model demonstrates significantly better memory retention than the Haar QRC baseline, with a long tail spanning beyond $d=50$. Importantly, this enhancement appears even though the convergence speed [quantified by $\rho_{A}^{(k)}-\rho_{B}^{(k)}||$ in Fig. S6(b)] is comparable to that observed at shorter $\Delta t_{\mathrm{in}}$. This discrepancy implies that, while the distance-based convergence metric is valuable for characterizing QRC systems, its direct connection to performance can be system-dependent. Specifically, in the SYK2 model, ESP is satisfied only asymptotically, and it remains nontrivial to determine how this asymptotic nature precisely affects performance. Further investigation, potentially employing alternative convergence metrics, is left for future work.
@@
Figures S8(c) and S8(d) extend Figs. S8(a) and S8(b), respectively, by presenting heatmaps of $R_{d}^{2}$ for $0\leq d\leq 20$ as the input interval $\Delta t_{\mathrm{in}}$ is varied. In the SYK4 model [Fig. S8(c)], a key observation is an oblique line of high performance that appears where $\Delta t_{\mathrm{in}}$ is shorter than the Thouless time. This sharp feature represents the performance enhancement at the temporal edge of many-body quantum chaos, consistent with the observation in Fig. 3(c) of the main text. Notably, as discussed there, the exact location of the performance peak is task-dependent. For the STM task investigated here, the peak shifts gradually to the nonchaotic side (smaller $\Delta t_{\mathrm{in}}$) as $d$ increases, leading to the observed oblique line. By contrast, a similar enhancement is absent in the SYK2 model [Fig. S8(d)]; instead, it shows a binary separation between a regime of sufficient memory retention at large $\Delta t_{\mathrm{in}}$ and poor retention at small $\Delta t_{\mathrm{in}}$, lacking any distinct performance peak.
@@
Figures S8(e) and S8(f) show heatmaps of the NMSE for the NARMA tasks using the SYK4 and SYK2 models, respectively. The vertical axis shows the NARMA order $n$, and the horizontal axis indicates the input interval $\Delta t_{\mathrm{in}}$. Consistent with the STM task results, a distinct performance peak is evident in the SYK4 model, whereas the SYK2 model shows only gradual performance changes. As discussed in the main text, the absence of performance peak in the SYK2 model indicates that the performance enhancement observed in the SYK4 model indeed originates from many-body quantum chaos. Notably, as the NARMA order $n$ increases, the peak broadens and eventually vanishes, illustrating a regime in which the edge-induced peak vanishes due to excessive task complexity.
@@ IV.2 Reservoir performance for the interpolated model
We additionally present performance heatmaps for the NARMA task in Fig. S9(e) for $\Delta t_{\mathrm{in}}=50$ and Fig. S9(f) for $\Delta t_{\mathrm{in}}=1$. The horizontal axis represents the coupling strength $\kappa_{2}/J_{4}$, and the vertical axis shows the NARMA order $n$. For small NARMA orders with a long input interval [Fig. S9(e)], the performance is distinctly enhanced around the chaotic boundary, corresponding to the parametric edge of many-body quantum chaos. This edge-induced performance enhancement is absent for the short input interval $\Delta t_{\mathrm{in}}=1$ in Fig. S9(f) because, as discussed in the main text, the QRC requires a sufficiently long input interval to manifest its quantum chaotic characteristics. As the NARMA order $n$ increases, the task becomes substantially more challenging, and the QRC can no longer represent it adequately, as evidenced by the degraded performance. In this regime with large $n$, the effective strategy is essentially limited to prioritizing the representation of contributions driven by distant past inputs, rather than capturing complex nonlinear details. Accordingly, in the high-$n$ regime the performance contours for the NARMA task [Fig. S9(e)] resemble those for the STM task [Fig. S9(c)].
IV.3 Nonlinear STM task
To further benchmark the QRC, we employ the nonlinear extension of the STM task [27, 34]. This task is designed to systematically vary the relative demands on memory and nonlinear processing. As in the standard STM task, a sequence of random inputs $\{u^{(k)}\}$, with $u^{(k)}\in[0,1]$, is provided to the quantum reservoir. For a given polynomial order $n$ and delay $d$, the target output at time step $k$ is defined as $\bar{y}^{(k)}=\left(u^{(k-d)}\right)^{n}$. The case $n=1$ recovers the standard (linear) STM task analyzed in the main text. By increasing $n$ at fixed $d$, the task selectively enhances the demand for nonlinear processing, while the memory requirement (determined by $d$) remains unchanged. The QRC performance on this task is quantified using the determination coefficient $R^{2}_{d}$.
We present the performance on the nonlinear STM task with delay $d=4$ around the temporal and parametric edges of many-body quantum chaos in Figs. S10(a) and S10(b), respectively. In particular, the temporal edge is studied using the SYK4 model, while the parametric edge is examined using the interpolated model at $\Delta t_{\mathrm{in}}=50$. In both cases, the performance decreases monotonically with increasing the nonlinearity order $n$, reflecting the greater task complexity. Notably, in the temporal domain, the distinct performance peak at the temporal edge persists from low- to high-order nonlinearities. The position of this peak appears to be independent of the nonlinearity order $n$, suggesting that the location of the temporal edge for this task is governed by the memory properties.
In the parametric domain, the performance peak is blurred for the linear STM task ($n=1$), because memory retention is strongest in the integrable regime. However, as the nonlinearity is introduced to the task via $n\geq 2$, QRC performance in the integrable regime degrades more rapidly than in the chaotic regime, reflecting the former’s limited nonlinear transformation capability. Consequently, a performance peak at the parametric edge of many-body quantum chaos becomes prominent for these tasks. Remarkably, with increasing $n$, the peak position shift towards smaller $\kappa_{2}/J_{4}$ values, i.e., closer to the chaotic regime. This shift reflects that operating closer to the chaotic side enhances the nonlinear transformation capacity, which is required to solve tasks with higher nonlinearity order $n$.
Figure S10: QRC performance on the nonlinear STM task with delay $d=4$ for orders $n=1$ to $5$. (a) Performance of the SYK4 model plotted as a function of the input interval $\Delta t_{\mathrm{in}}$. The Thouless time $t_{\mathrm{th}}$ is indicated by the vertical dashed line. (b) Performance of the interpolated model as a function of the coupling strength $\kappa_{2}/J_{4}$, with the input interval fixed at $\Delta t_{\mathrm{in}}=50$. The onsets of the chaotic and integrable regimes are marked by the left and right horizontal lines, respectively.
IV.4 PC task
We next evaluate the PC task [17, 23, 29, 34]. In this task, binary inputs ${u^{\prime}}^{(k)}\in\{0,1\}$ are injected into the quantum reservoir, and the target output is the parity of the sum of the previous inputs: $\bar{y}^{(k)}=\left(\sum_{d=0}^{n}u^{\prime(k-d)}\right)\bmod 2$. We also use the determination coefficient $R^{2}$ to evaluate performance on this task. A defining feature of this task is its strong nonlinearity requirement. The PC task of order $n$ requires memory spanning $n+1$ steps (the current input plus the $n$ preceding ones). More importantly, it requires nonlinearity of order $n+1$ involving many cross-term interactions among delayed inputs, because the parity function corresponds to an ($n+1$)-degree polynomial in its inputs. In other words, increasing $n$ by one adds only a single additional step to the memory requirement, whereas the nonlinearity requirement grows much more sharply both in polynomial order and in the combinatorial complexity of cross-term interactions.
Figure S11: QRC performance on (a), (b) the PC task and (c), (d) the PC task with memory bias $b=2$. (a), (c) Performance using the SYK4 model as a function of $\Delta t_{\mathrm{in}}$. (b), (d) Performance using the interpolated model as a function of $\kappa_{2}/J_{4}$, with $\Delta t_{\mathrm{in}}=50$. In all panels, $R^{2}$ values for $n=1$ are plotted on the left vertical axis, and values for $n=2$ and 33 are on the right vertical axis.
Figures S11(a) and S11(b) presents the QRC performance for the PC task with $n=1,2$, and $3$, evaluated in the temporal domain [Fig. S11(a)] and the parametric domain [Fig. S11(b)]. For $n=1$, we observe clear performance peaks at the edge of many-body quantum chaos in both cases. In contrast, for $n=2$ and $3$, the task’s nonlinearity requirement escalates sharply and overwhelms the memory demand. Consequently, optimal performance is achieved within the chaotic regime, where the nonlinearity of the input-to-output mapping is strongest. This behavior exemplifies an extreme imbalance in task requirements: when either memory or nonlinear transformation dominates, the optimal operating regime is integrable or chaotic, respectively, and the edge-induced performance peak is absent. The linear STM task (memory-dominated) and the PC task with $n\geq 2$ (nonlinearity-dominated) exemplify these extremes. It should be noted that in such limits, explicit design guidelines are less necessary because the optimal operating regime follows directly from the task definition.
Having observed that the edge-induced peak appears when a nonlinearity requirement is added to an otherwise memory-dominant STM task (the nonlinear STM task in Fig. S10), we next examine whether introducing an additional memory requirement to the PC task likewise produces such a peak. To this end, we modify the target output to $\bar{y}^{(k)}=\left(\sum_{d=0}^{n}u^{\prime(k-d-b)}\right)\bmod 2$, where $b$ denotes the temporal bias step. Figures S11(c) and S11(d) present the QRC performance for the PC task $n=1,2,3$ under the bias $b=2$. Strikingly, a clear performance enhancement at the edge of many-body quantum chaos is observed for both edges, across all orders investigated. Because the bias $b$ increases only the memory requirement while leaving the nonlinearity unchanged, these results support the interpretation that the absence of a performance peak in the original PC task is due to the dominance of the nonlinear requirement. Introducing the bias makes memory a performance limiting factor alongside nonlinearity, thereby resolving the imbalance in task demands. The resulting setting thus becomes analogous to the NARMA task and the nonlinear STM task, for which both memory and nonlinearity constrain performance; in such nontrivial regimes, the edge of many-body quantum chaos becomes most prominent. Although not known a priori, realistic scenarios involving real-world data are expected to fall into this category. The pronounced performance peaks observed for these tasks therefore highlight the importance of operating at the edge of many-body quantum chaos as a robust design guideline for QRC. Finally, we note that increasing the order $n$ in the biased PC task shifts the peak toward the chaotic regime, paralleling the trend observed in the nonlinear STM task (Fig. S10).
IV.5 Impact of statistical uncertainty in measurements
Figure S12: The QRC performance (NMSE) on the NARMA-5 task under sampling noise. (a) The NMSE for the SYK4 model as a function of the input interval $\Delta t_{\mathrm{in}}$. (b) The NMSE for the interpolated model as a function of the coupling strength $\kappa_{2}/J_{4}$, with $\Delta t_{\mathrm{in}}$ fixed at $50$. In both panels, the colors and markers indicate the sampling noise strength $\sigma$.
Because expectation values are estimated from a finite number of measurement shots, their accuracy is limited by sampling noise. Our analysis thus far has assumed ideal, noise-free conditions; we now investigate the impact of statistical noise on QRC performance by modeling finite sampling error as additive Gaussian noise. Specifically, to the noiseless expectation value $n_{i}=\langle c_{i}^{\dagger}c_{i}\rangle$, we add a zero-mean Gaussian noise with standard deviation $\sigma_{i}=\sigma\sqrt{n_{i}(1-n_{i})}$, where $\sigma$ parametrizes the shot noise strength [26, 30, 32]. This variance matches the binomial variance of repeated projective measurements in the $c_{i}^{\dagger}c_{i}$ basis, and the Gaussian approximation is valid when $\sigma\ll 1$.
In Figs. S12(a) and S12(b), we show the NARMA-5 performance around the temporal edge and parametric edge, respectively, as the sampling noise strength $\sigma$ is varied. As expected for algorithms that rely on estimating expectation values, statistical fluctuations degrade the performance; accordingly, reducing $\sigma$ monotonically improves it. Note that the sensitivity to this noise depends on the operating regime. In the temporal sweep [Fig. S12(a)], small input intervals $\Delta t_{\mathrm{in}}$ are markedly more vulnerable than large $\Delta t_{\mathrm{in}}$ (the RMT regime), because the system-specific structure is more pronounced at small $\Delta t_{\mathrm{in}}$ and is therefore more susceptible to statistical noise. For large $\sigma$, this degradation overwhelms the edge-induced enhancement, producing local minima near the temporal edge and shifting the global minimum to the large-$\Delta t_{\mathrm{in}}$ regime. As $\sigma$ is reduced to $10^{-8}$, the temporal edge again yields the global minimum; at $\sigma=10^{-12}$, it nearly recovers the ideal performance shown in Fig. 3(c) of the main text. By contrast, in the parametric sweep we fix $\Delta t_{\mathrm{in}}=50$, where the QRC is comparatively robust to sampling noise. Consequently, the parametric edge of many-body quantum chaos attains the global minimum even at $\sigma=10^{-3}$, and the performance is essentially converged to the ideal from $\sigma=10^{-4}$ onward.